I am not sure this is the right area, we are studying real number exponents.
6^sqrt2 * 6^sqrt8/6^3sqrt2 simplify
I hope it makes sense, here is long hand.
6 to the square root of 2 times 6 to the square root of 8 all over (or divided by) 6 to the power 3 to the square root of 2.
except for the sixes everything else is a power.

Answer by stanbon(75887)

[6^sqrt2 * 6^sqrt8]/[6^3sqrt2] simplify
---------------------------------
= 6^(sqrt(2)+sqrt(8)]/[6^3sqrt(2]
--------------------
= 6^(sqrt(8)-2sqrt(2)]
----
= 6^(2sqrt(2)-2sqrt(2)]
----
= 6^0
-----
= 1
=============
Cheers,
Stan H.